OpenStudy (anonymous):

Find the following limit lim 3xcsc(2pi-x) x->0+

OpenStudy (anonymous):

usual gimmick is to write it as \[3\lim_{x\to 0^+}\frac{\csc(2\pi-x)}{\frac{1}{x}}\] and use l'hopital's rule

OpenStudy (anonymous):

So use product rule and go from there?

OpenStudy (anonymous):

oh no take the derivative top and bottom separately

OpenStudy (anonymous):

top is -cotxcscx(2pi-x)(-1) = cotxcscx(2pi-x) bottom is -1/x^2 Correct?